# A Survey of Long-Document Retrieval in the PLM and LLM Era

### v1

@@ Abstract.
The proliferation of long-form documents presents a fundamental challenge to information retrieval (IR), as their length, dispersed evidence, and complex structures demand specialized methods beyond standard passage-level techniques. This survey provides the first comprehensive treatment of long-document retrieval (LDR), consolidating methods, challenges, and applications across three major eras. We systematize the evolution from classical lexical and early neural models to modern pre-trained (PLM) and large language models (LLMs), covering key paradigms like passage aggregation, hierarchical encoding, efficient attention, and the latest LLM-driven re-ranking and retrieval techniques. Beyond the models, we review domain-specific applications, specialized evaluation resources, and outline critical open challenges such as efficiency trade-offs, multimodal alignment, and faithfulness. This survey aims to provide both a consolidated reference and a forward-looking agenda for advancing long-document retrieval in the era of foundation models.
@@ 1. Introduction and Contribution
From locating a single critical clause in a multi-hundred-page legal contract to synthesizing evidence from thousands of biomedical research papers, the ability to retrieve precise information from long-form documents has become a foundational challenge in the modern information landscape. The exponential growth of digital information has yielded corpora where such documents—scientific articles and patents, statutes and judicial opinions, financial and technical reports, clinical notes, books, and multimedia-rich webpages—are the primary carriers of knowledge. These artifacts routinely span thousands to tens of thousands of tokens and exhibit rich internal structure. Retrieving evidence from such materials are foundational to web search, legal discovery, scientific knowledge mining, clinical decision support, and enterprise intelligence. Yet the very properties that make long documents valuable—distributed evidence, hierarchical organization, and inter-document linkage—also make them challenging for conventional information retrieval pipelines.
@@
Building on these foundations, the LLM era has introduced instruction-following models for two primary roles: as powerful re-rankers, exemplified by listwise prompting approaches like RankGPT, and—through fine-tuning—as end-to-end retrievers, such as the bi-encoder variants RepLLaMA and RankLLaMA. This new wave is also accompanied by system-level innovations for long-context efficiency, including sparse attention for LLMs, prompt compression, and KV-cache reuse.
@@
To the best of our knowledge, this is the first survey to offer a holistic and consolidated view specifically on long-document retrieval, tracing its progression across three distinct eras—from classical lexical models to the latest PLM and LLM-based paradigms. Our work distinguishes itself by not only systematizing the models but also providing integrated guidance on domain-specific applications, evaluation benchmarks, and a forward-looking research agenda, thus offering a complete reference for the field.
@@
  • Efficiency principles for long contexts. We distill design patterns—sparse attention, hierarchical pooling, query-focused routing, prompt compression, and KV-cache reuse—and relate them to first-stage recall, re-ranking, and reading/generation.
  • Evaluation guidance. We consolidate text-only, structure-aware, multimodal, and cross-lingual benchmarks; discuss segment-level metrics, partial pooling, and reliability under sparse judgments; and outline best practices for reproducible large-corpus evaluation.
@@
  • A forward-looking research agenda. We articulate critical gaps in the field, including label sparsity, robust long-query handling, faithful multimodal retrieval, and efficiency–effectiveness trade-offs, outlining promising directions for future work.
Roadmap. The survey first reviews pre-PLM baselines, then covers PLM-based LDR and LLM-based retrievers/rerankers with efficiency innovations and multimodal extensions. We next provide a comparative analysis, summarize datasets and evaluation protocols—including structure- and layout-aware benchmarks—and detail application blueprints in law, biomedicine, web/news, scientific retrieval, and cross-lingual settings. We conclude with open challenges and a research agenda.
@@
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=10.5em,fill=fill_0,draw=draw_0,font=,, where level=1text width=11.5em,fill=fill_1,draw=draw_1,font=,, where level=2text width=14em,fill=fill_2,draw=draw_2,font=,, where level=3font=,, [ Long-Document
Retrieval, minimum height=2.8em [ Pre-PLM Era
(Sec 3) [ Lexical Methods
(Sec 3.1) [ TF-IDF (Jones, 1973), BM25 (Robertson et al., 1995), Heuristic Chunking (Callan, 1994), leaf, text width=42em ] ] [ Early Neural IR
(Sec 3.2) [ DSSM (Huang et al., 2013), DRMM (Guo et al., 2016), MatchPyramid (Pang et al., 2016), PACRR (Hui et al., 2017), DeepRank (Pang et al., 2017), leaf, text width=42em ] ] ] [ Core Paradigms
PLM & LLM Era
(Sec 4) [ The Holistic Paradigm
(Sec 4.1) [ Naive Truncation: DPR (Karpukhin et al., 2020), ANCE (Xiong et al., 2020), leaf, text width=42em ] [ Long-Sequence Transformers: Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020b), Reformer (Kitaev et al., [n. d.]), FlashAttention (Dao et al., 2022),
LongNet (Ding et al., 2023), LONGEMBED (Zhu et al., 2024), M3-Embedding (Chen et al., 2024), QDS-Transformer (Jiang et al., 2020), Socialformer (Zhou et al., 2022), leaf, text width=42em ] [ LLMs as Rerankers: RankGPT (Sun et al., 2023), RepLLaMA/RankLLaMA (Ma et al., 2024), leaf, text width=42em ] [ Efficiency Mitigation: LongLoRA (Chen et al., 2023), LongLLMLingua (Jiang et al., 2024), DuoAttention (Xiao et al., [n. d.]), leaf, text width=42em ] ] [ The Divide-and-Conquer
Paradigm (Sec 4.2) [ Pooling-based Heuristics: BERT-MaxP/SumP (Dai and Callan, 2019), leaf, text width=42em ] [ Hierarchical Aggregation: PARADE (Li et al., 2023b), DRSCM (Wang et al., 2023), LTR-BERT (Wang et al., 2024), MORES+ (Gao and Callan, 2022), leaf, text width=42em ] [ Key Block Selection: IDCM (Hofstätter et al., 2021), KeyB (Li et al., 2023a), KeyB2 (Li et al., 2024), DCS (Sheng et al., 2025), leaf, text width=42em ] [ Hybrid & Late Interaction: ICLI (Li and Gaussier, 2022), Match-Ignition (Pang et al., 2021a), Longtriever (Yang et al., 2023), leaf, text width=42em ] ] [ Indexing-Structure-
Oriented Paradigm
(Sec 4.3) [ MC-indexing (Dong et al., 2024), HELD (Cao et al., 2022), RAPTOR (Sarthi et al., 2024), leaf, text width=42em ] ] [ Long-Query Retrieval
(Sec 4.4) [ RPRS (Askari et al., 2024), leaf, text width=42em ] ] ] [ Evaluation
(Sec 5) [ Datasets and Benchmarks
(Sec 5.1) [ TREC Robust04 (Voorhees, 2004), Gov2/Terabyte Track (Büttcher et al., 2006), ClueWeb12 (Lemur Project, 2012), TREC Deep Learning (Craswell et al., 2020), MS MARCO (Bajaj et al., 2016),
MLDR  (Chen et al., 2024), LONGEMBED (Zhu et al., 2024), Multi-CPR (Chinese) (Long et al., 2022), leaf, text width=42em ] ] [ Standard Metrics
(Sec 5.2) [ nDCG@k, MAP, Recall@k, Precision@k, MRR,ERR, leaf, text width=42em ] ] ] [ Applications
(Sec 6) [ Legal Retrieval
(Sec 6.1) [ Lawformer (Xiao et al., 2021), LegalBERT (Chalkidis et al., 2020), LawGPT (Zhou et al., 2024), ChatLaw (Cui et al., 2023), DISC-LawLLM (Yue et al., 2023), leaf, text width=42em ] ] [ Scholarly Paper Retrieval
(Sec 6.2) [ SciBERT (Beltagy et al., 2019a), SPECTER (Cohan et al., 2020), SciNCL (Ostendorff et al., 2022), PRISM (Park et al., 2025), CORANK (Tian et al., 2025), PaperQA (Lála et al., 2023), SemRank (Zhang et al., 2025),
DocReLM (Wei et al., 2024), LLM-KnowSimFuser (Dai et al., 2024), PaSa (He et al., 2025), SPAR (Shi et al., 2025a), leaf, text width=42em ] ] [ Biomedical Retrieval
(Sec 6.3) [ PubMedBERT (Gu et al., 2021), Clinical-Longformer, Clinical-BigBird (Li et al., 2022), BioClinical ModernBERT (Sounack et al., 2025), leaf, text width=42em ] ] [ Cross-Lingual Retrieval
(Sec 6.4) [ mBERT (Devlin et al., 2018), LaBSE (Feng et al., 2020), XLM-R, McCrolin (Limkonchotiwat et al., 2024), mGTE (Zhang et al., 2024), CROSS (Nezhad and Agrawal, 2025), leaf, text width=42em ] ] ] ]
@@ 2. Problem Scope And Definition
Long-Document Retrieval is an important branch task in information retrieval. Its goal is to retrieve the most relevant documents or document fragments based on the user’s query when facing a corpus containing a large number of long documents. Compared with classic document retrieval or paragraph retrieval tasks, LDR is unique in that it not only requires accurate identification of relevant information in a large document, but also requires efficient processing of redundant information and structural complexity while maintaining semantic integrity. As illustrated in Fig. 1 , a clear distinction emerges between two types of retrieval methods when the query target is information within long documents. General retrieval methods struggle to capture scattered and detailed information completely, making it difficult to extract key content from long documents and present it effectively. In contrast, long-document retrieval methods are specifically designed to address this issue. Their core strength lies in the ability to accurately retrieve deep information from large-scale textual resources and organize such information in a structured manner, thereby fulfilling query requirements in long-document scenarios. The figure visually demonstrates the irreplaceability of long-document retrieval methods in specific contexts and provides clear guidance for the design of future methods. This capability gap also explains why ”long documents” need to be defined from multiple dimensions (length, structure, and semantics) —— these characteristics are exactly the root causes of the ineffectiveness of general retrieval methods.
@@ 2.1.1. Sparse and Inconsistent Supervision Signals
In the majority of public benchmarks, relevance annotations for lengthy documents are typically concentrated at the document or paragraph level, while actual queries frequently involve only small fragments within the document. As a result of this discrepancy, the model’s discriminative power is diminished, and fine-grained alignment becomes more difficult. For instance, (Kwiatkowski et al., 2019)’s Natural Questions dataset offers paragraph-level ’long answers’ (typically a paragraph) and brief answer annotations for each question, which somewhat parallel the ’query–local fragment’ correspondence. Nevertheless, this annotation type continues to fall short of addressing the fine-grained supervisory requirements at the sentence or even evidence chain level. Conversely, sparse relevance judgements are typically implemented by mainstream training and evaluation resources in the retrieval domain, including MS MARCO and TREC Deep Learning. The official MS MARCO leaderboard is based on ”pre-collected sparse judgements”, as explicitly stated by (Craswell et al., 2020) in their TREC 2022 review. Additionally, the sparse character of MS MARCO labels is underscored in the TREC DL overview. (Arabzadeh et al., 2022) conducted an analysis of this sparse supervision method, observing that it may result in external validity issues in evaluations by obstructing the learning of more detailed matching patterns by models.
@@ 2.1.2. Document Segmentation and Chunking
To facilitate indexing and retrieval, it is frequently necessary to divide lengthy documents into smaller blocks, such as paragraphs or chunks. A long-standing challenge is determining the optimal segmentation without sacrificing semantic context. As a result of mechanical segmentation, traditional fixed-length chunking frequently disrupts sentence logic or conceptual units, resulting in ”semantic fragmentation” during retrieval. Through multi-dataset analysis, (Bhat et al., 2025) systematically evaluated the influence of fixed-length chunking strategies on retrieval performance. They discovered that smaller chunks (64–128 tokens) are more effective in short-answer scenarios, as they are able to accurately identify individual factual information. Conversely, larger chunks (512–1024 tokens) are more advantageous for tasks that necessitate a broader contextual understanding, such as thematic summarisation of long documents and novel plot analysis. Different embedding models exhibit varying sensitivities to block size, which they emphasise. For example, lightweight embedding models are more efficient when using smaller chunks, whereas large parameter models necessitate larger chunks to completely capture deep semantic information. This renders the selection of a segmentation strategy a complex yet essential step in the pursuit of a balance between semantic integrity and retrieval efficiency.
@@ 2.1.3. Computational and Efficiency Issues
The computational and storage overhead are substantially increased when long texts are processed. The retrieval throughput and latency are influenced by the larger index sizes and the increased number of word vectors or embeddings that must be computed when the documents are longer. This problem is especially evident in deep learning-based retrieval, particularly in Transformer-based re-ranking models. The computational complexity of the self-attention mechanism increases quadratically with the length of the input, rendering it nearly impossible to process ultra-long sequences in a single pass. Researchers have suggested a variety of enhancements in various directions to resolve this bottleneck. On the one hand, BigBird (Zaheer et al., 2020b) and Longformer (Beltagy et al., 2020) employ sparse attention mechanisms to effectively enhance the model’s capacity to manage lengthy sequences, thereby reducing computational complexity at the structural level. In contrast, (Seo et al., 2024) proposed Colour (Compression for Long Context Retrieval), an input-level approach that reduces input size by compressing paragraphs and retaining core information, thereby avoiding redundant computations caused by irrelevant content. The former enhances scalability by optimising the model structure, whereas the latter achieves cost reduction and efficiency gains through content compression.

### v2

@@ Abstract.
The rapid growth of long-form documents presents a fundamental challenge to information retrieval (IR). Their length, dispersed evidence, and complex structures demand approaches that go beyond standard passage-level techniques. This survey provides a comprehensive survey of long-document retrieval (LDR) and organizes existing work into three major eras of development. The survey traces progress from classical lexical and early neural models to pre-trained language models (PLMs) and large language models (LLMs). It highlights core paradigms such as passage aggregation, hierarchical encoding, efficient attention, and LLM-driven re-ranking and retrieval. In addition to methods, the survey summarizes domain-specific applications, available evaluation resources, and the benchmarks that shape empirical study. It identifies open challenges related to efficiency trade-offs, evidence localization, multimodal alignment, and interpretability. The survey primarily focuses on methodological advances, while also considering practical applications in diverse domains. The principal conclusion is that future progress in LDR will depend on hybrid solutions that combine sub-linear indexing, structure-aware modeling, and LLM-based reasoning, as identified through the methods and challenges systematized in this survey.
@@ 1. Introduction and Contribution
From locating a single critical clause in a multi-hundred-page legal contract to synthesizing evidence from thousands of biomedical research papers, the ability to retrieve precise information from long-form documents has become a foundational challenge in the modern information landscape. The exponential growth of digital information has yielded corpora where such documents-scientific articles and patents, statutes and judicial opinions, financial and technical reports, clinical notes, books, and multimedia-rich web pages are the primary carriers of knowledge. These artifacts routinely span thousands to tens of thousands of tokens and exhibit rich internal structure. Retrieving evidence from such materials are foundational to web search, legal discovery, scientific knowledge mining, clinical decision support, and enterprise intelligence. Yet the very properties that make long documents valuable, such as distributed evidence, hierarchical organization, and inter-document linkage, also make them challenging for conventional information retrieval pipelines.
@@
Building on these foundations, the LLM era has introduced instruction-following models for two primary roles: as powerful re-rankers, exemplified by listwise prompting approaches like RankGPT, and through fine-tuning as end-to-end retrievers, such as the bi-encoder variants RepLLaMA and RankLLaMA. This new wave is also accompanied by system-level innovations for long-context efficiency, including sparse attention for LLMs, prompt compression, and KV-cache reuse.
@@
To the best of our knowledge, this is the first survey to offer a holistic and consolidated view specifically on long-document retrieval, tracing its progression across three distinct eras: from classical lexical models to the latest PLM and LLM-based paradigms. Our work distinguishes itself by not only systematizing the models but also providing integrated guidance on domain-specific applications, evaluation benchmarks, and a forward-looking research agenda, thus offering a complete reference for the field.
@@
  • Efficiency principles for long contexts. We distill design patterns: sparse attention, hierarchical pooling, query-focused routing, prompt compression, KV-cache reuse, and relate them to first-stage recall, re-ranking, and reading/generation.
@@
  • A forward-looking research agenda. We articulate critical gaps in the field, including label sparsity, robust long-query handling, faithful multimodal retrieval, and efficiency-effectiveness trade-offs, outlining promising directions for future work.
Roadmap. The survey first reviews pre-PLM baselines, then covers PLM-based LDR and LLM-based retrievers/rerankers with efficiency innovations and multimodal extensions. We next provide a comparative analysis, summarize datasets and evaluation protocols, and detail application blueprints in law, biomedicine, web/news, scientific retrieval, and cross-lingual settings. We conclude with open challenges and a research agenda.
@@
Long-Document Retrieval Pre-PLM Era (Sec 3) Lexical Methods (Sec 3.1) TF-IDF (Jones, 1973), BM25 (Robertson et al., 1995), Heuristic Chunking (Callan, 1994) Early Neural IR (Sec 3.2) DSSM (Huang et al., 2013), DRMM (Guo et al., 2016), MatchPyramid (Pang et al., 2016), PACRR (Hui et al., 2017), DeepRank (Pang et al., 2017) Core Paradigms PLM & LLM Era (Sec 4) The Holistic Paradigm (Sec 4.1) Naive Truncation: DPR (Karpukhin et al., 2020), ANCE (Xiong et al., 2020) Long-Sequence Transformers: Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020b), Reformer (Kitaev et al., [n. d.]), FlashAttention (Dao et al., 2022), LongNet (Ding et al., 2023), LONGEMBED (Zhu et al., 2024), M3-Embedding (Chen et al., 2024), QDS-Transformer (Jiang et al., 2020), Socialformer (Zhou et al., 2022) LLMs as Rerankers: RankGPT (Sun et al., 2023), RepLLaMA/RankLLaMA (Ma et al., 2024) Efficiency Mitigation: LongLoRA (Chen et al., 2023), LongLLMLingua (Jiang et al., 2024), DuoAttention (Xiao et al., [n. d.]) The Divide-and-Conquer Paradigm (Sec 4.2) Pooling-based Heuristics: BERT-MaxP/SumP (Dai and Callan, 2019) Hierarchical Aggregation: PARADE (Li et al., 2023b), DRSCM (Wang et al., 2023), LTR-BERT (Wang et al., 2024), MORES+ (Gao and Callan, 2022) Key Block Selection: IDCM (Hofstätter et al., 2021), KeyB (Li et al., 2023a), KeyB2 (Li et al., 2024), BReps (Li et al., 2025), DCS (Sheng et al., 2025) Hybrid & Late Interaction: ICLI (Li and Gaussier, 2022), Match-Ignition (Pang et al., 2021a), Longtriever (Yang et al., 2023) Indexing-Structure- Oriented Paradigm (Sec 4.3) MC-indexing (Dong et al., 2024), HELD (Cao et al., 2022), RAPTOR (Sarthi et al., 2024) Long-Query Retrieval (Sec 4.4) RPRS (Askari et al., 2024) Evaluation (Sec 5) Datasets and Benchmarks (Sec 5.1) TREC Robust04 (Voorhees, 2004), Gov2/Terabyte Track (Büttcher et al., 2006), ClueWeb12 (Lemur Project, 2012), TREC Deep Learning (Craswell et al., 2020), MS MARCO (Bajaj et al., 2016), MLDR  (Chen et al., 2024), LONGEMBED (Zhu et al., 2024), Multi-CPR (Chinese) (Long et al., 2022) Standard Metrics (Sec 5.2) nDCG@k, MAP, Recall@k, Precision@k, MRR,ERR Applications (Sec 6) Legal Retrieval (Sec 6.1) Lawformer (Xiao et al., 2021), LegalBERT (Chalkidis et al., 2020), LawGPT (Zhou et al., 2024), ChatLaw (Cui et al., 2023), DISC-LawLLM (Yue et al., 2023) Scholarly Paper Retrieval (Sec 6.2) SciBERT (Beltagy et al., 2019a), SPECTER (Cohan et al., 2020), SciNCL (Ostendorff et al., 2022), PRISM (Park et al., 2025), CORANK (Tian et al., 2025), PaperQA (Lála et al., 2023), SemRank (Zhang et al., 2025), DocReLM (Wei et al., 2024), LLM-KnowSimFuser (Dai et al., 2024), PaSa (He et al., 2025), SPAR (Shi et al., 2025a) Biomedical Retrieval (Sec 6.3) PubMedBERT (Gu et al., 2021), Clinical-Longformer, Clinical-BigBird (Li et al., 2022), BioClinical ModernBERT (Sounack et al., 2025) Cross-Lingual Retrieval (Sec 6.4) mBERT (Devlin et al., 2018), LaBSE (Feng et al., 2020), XLM-R, McCrolin (Limkonchotiwat et al., 2024), mGTE (Zhang et al., 2024), CROSS (Nezhad and Agrawal, 2025)
@@ 2. Problem Scope And Definition
Long-Document Retrieval is an important branch task in information retrieval. Its goal is to retrieve the most relevant documents or document fragments based on the user’s query when facing a corpus containing a large number of long documents. Compared with classic document retrieval or paragraph retrieval tasks, LDR is unique in that it not only requires accurate identification of relevant information in a large document, but also requires efficient processing of redundant information and structural complexity while maintaining semantic integrity. As illustrated in Fig. 1 , a clear distinction emerges between two types of retrieval methods when the query target is information within long documents. General retrieval methods struggle to capture scattered and detailed information completely, making it difficult to extract key content from long documents and present it effectively. In contrast, long-document retrieval methods are specifically designed to address this issue. Their core strength lies in the ability to accurately retrieve deep information from large-scale textual resources and organize such information in a structured manner, thereby fulfilling query requirements in long-document scenarios. The figure visually demonstrates the irreplaceability of long-document retrieval methods in specific contexts and provides clear guidance for the design of future methods. This capability gap also explains why ”long documents” need to be defined from multiple dimensions (length, structure, and semantics). These characteristics are exactly the root causes of the ineffectiveness of general retrieval methods.
@@ 2.1.1. Sparse and Inconsistent Supervision Signals
In the majority of public benchmarks, relevance annotations for lengthy documents are typically concentrated at the document or paragraph level, while actual queries frequently involve only small fragments within the document. As a result of this discrepancy, the model’s discriminative power is diminished, and fine-grained alignment becomes more difficult. For instance, the Natural Questions dataset introduced by Kwiatkowski(Kwiatkowski et al., 2019) provides paragraph-level long answers and brief answer annotations for each question, which somewhat parallel the query-local fragment correspondence. Nevertheless, this annotation type continues to fall short of addressing the fine-grained supervisory requirements at the sentence or even evidence chain level. Conversely, sparse relevance judgements are typically implemented by mainstream training and evaluation resources in the retrieval domain, including MS MARCO and TREC Deep Learning. The official MS MARCO leaderboard is based on pre-collected sparse judgements, as explicitly stated by Craswell(Craswell et al., 2020) in their TREC 2022 review. Additionally, the sparse character of MS MARCO labels is underscored in the TREC DL overview. Arabzadeh(Arabzadeh et al., 2022) conducted an analysis of this sparse supervision method, observing that it may result in external validity issues in evaluations by obstructing the learning of more detailed matching patterns by models.
@@ 2.1.2. Document Segmentation and Chunking
To facilitate indexing and retrieval, it is frequently necessary to divide lengthy documents into smaller blocks, such as paragraphs or chunks. A long-standing challenge is determining the optimal segmentation without sacrificing semantic context. As a result of mechanical segmentation, traditional fixed-length chunking frequently disrupts sentence logic or conceptual units, resulting in ”semantic fragmentation” during retrieval. Through multi-dataset analysis, Bhat(Bhat et al., 2025) systematically evaluated the influence of fixed-length chunking strategies on retrieval performance. They discovered that smaller chunks are more effective in short-answer scenarios, as they are able to accurately identify individual factual information. Conversely, larger chunks are more advantageous for tasks that necessitate a broader contextual understanding, such as thematic summarisation of long documents and novel plot analysis. Different embedding models exhibit varying sensitivities to block size, which they emphasise. For example, lightweight embedding models are more efficient when using smaller chunks, whereas large parameter models necessitate larger chunks to completely capture deep semantic information. This renders the selection of a segmentation strategy a complex yet essential step in the pursuit of a balance between semantic integrity and retrieval efficiency.
@@ 2.1.3. Computational and Efficiency Issues
The computational and storage overhead are substantially increased when long texts are processed. The retrieval throughput and latency are influenced by the larger index sizes and the increased number of word vectors or embeddings that must be computed when the documents are longer. This problem is especially evident in deep learning-based retrieval, particularly in Transformer-based re-ranking models. The computational complexity of the self-attention mechanism increases quadratically with the length of the input, rendering it nearly impossible to process ultra-long sequences in a single pass. Researchers have suggested a variety of enhancements in various directions to resolve this bottleneck. On the one hand, BigBird (Zaheer et al., 2020b) and Longformer (Beltagy et al., 2020) employ sparse attention mechanisms to effectively enhance the model’s capacity to manage lengthy sequences, thereby reducing computational complexity at the structural level. In contrast, Seo(Seo et al., 2024) proposed Colour (Compression for Long Context Retrieval), an input-level approach that reduces input size by compressing paragraphs and retaining core information, thereby avoiding redundant computations caused by irrelevant content. The former enhances scalability by optimising the model structure, whereas the latter achieves cost reduction and efficiency gains through content compression.